# Aligning Individual and Collective Objectives in Multi-Agent Cooperation

### v1

@@ Abstract
In the field of multi-agent learning, the challenge of mixed-motive cooperation is pronounced, given the inherent contradictions between individual and collective goals. Current research in this domain primarily focuses on incorporating domain knowledge into rewards or introducing additional mechanisms to foster cooperation. However, many of these methods suffer from the drawbacks of manual design costs and the lack of a theoretical grounding convergence procedure to the solution. To address this gap, we approach the mixed-motive game by modeling it as a differentiable game to study learning dynamics. We introduce a novel optimization method named Altruistic Gradient Adjustment (AgA) that employs gradient adjustments to novelly align individual and collective objectives. Furthermore, we provide theoretical proof that the selection of an appropriate alignment weight in AgA can accelerate convergence towards the desired solutions while effectively avoiding the undesired ones. The visualization of learning dynamics effectively demonstrates that AgA successfully achieves alignment between individual and collective objectives. Additionally, through evaluations conducted on established mixed-motive benchmarks such as the public good game, Cleanup, Harvest, and our modified mixed-motive SMAC environment, we validate AgA’s capability to facilitate altruistic and fair collaboration.
Machine Learning, ICML
[…]
@@ 1 Introduction
Multi-agent cooperation primarily focuses on learning how to promote collaborative behavior in shared environments. Typically, research divides multi-agent cooperation into two prominent areas, pure-motive cooperation and mixed-motive cooperation (Du et al., 2023; McKee et al., 2020a). Recent progress in cooperative Multi-Agent Reinforcement Learning (MARL) has primarily focused on addressing pure-motive cooperation, which also known as a common-payoff game, represents situations where individual goals fully align with collective objectives (Yu et al., 2022; Wang et al., 2020; Zhong et al., 2023; Li et al., 2024). Despite pure-motive cooperation, mixed-motive cooperation, which is typically defined by the imperfect alignment between individual and collective rationalities, is more common in real-world scenarios (Rapoport, 1974; McKee et al., 2020a).
@@
Recent mixed-motive cooperative studies in MARL have largely employed hand-crafted designs to promote collaboration. Incorporating additional mechanisms to align objectives appears to be a popular avenue, including aspects such as reputation (Anastassacos et al., 2021), norms (Vinitsky et al., 2023), and contracts (Hughes et al., 2020). Another prevalent method draws on intrinsic motivation to align individual and collective objectives, augmenting altruistic collaboration by integrating heuristic knowledge into the incentive function. Conventionally, some works accumulate individual rewards with the group to cultivate altruistic conduct (Hostallero et al., 2020; Peysakhovich & Lerer, 2018; Apt & Schäfer, 2014). Furthermore, some studies derive more sophisticated preference signals from the rewards of other agents (Hughes et al., 2018; McKee et al., 2020a). Additionally, several approaches aim to learn the potential influences of an agent’s actions on others (Yang et al., 2020; Jaques et al., 2019; Lu et al., 2022). Most of these algorithms are highly dependent on carefully crafted designs, necessitating significant human expertise and detailed domain knowledge. However, they lack in-depth theoretical analysis concerning alignment and convergence.
In this work, we first present the differentiable mixed-motive game (DMG), premised on the assumption that the players’ loss functions are at least twice differentiable. The proposed formulation offers an efficient tool for analyzing the dynamics of gradient-based methods when optimizing mixed-motive objectives. Furthermore, we explore the optimization trajectories of simultaneous optimization methods and gradient adjustment methods ( consensus optimization (Mescheder et al., 2017) and symplectic gradient adjustment (Letcher et al., 2019)) in the context of a two-player DMG, as shown in Fig. 1. First, we observe that gradient adjustment methods are ineffective in addressing mixed-motive problems. These methods tend to prioritize stability at the expense of mitigating individual loss, which conflicts with the interests of individual participants (Chen et al., 2023). While simultaneous optimization with collective loss successfully enhances collective interests, it seems to neglect individual interests.
To address the issue, we propose an Altruistic Gradient Adjustment (AgA) optimization algorithm, designed to align individual and collective objectives from a gradient perspective. We provide a theoretical proof to suggest that appropriately choosing the sign of the alignment weight of the gradient terms could speed up the convergence of the AgA optimization strategy towards stable fixed point of collective objective while effectively bypassing unstable ones. The trajectory visualizations in Fig. 1, along with the sign selection ablation study trajectories presented in Fig. 3, serve to demonstrate successful alignment of objectives.
To further assess the effectiveness of our proposed AgA approach, we perform an assortment of experiments on a two-player public goods game and two sequential social dilemma games (Cleanup and Harvest) (Leibo et al., 2017). Given the relatively simple nature of these standard test scenarios used in mixed-motive cooperation contexts, we chose to modify the widely used SMAC environment (Samvelyan et al., 2019) to create a more complex mixed-motive cooperative setting. Evaluation based on various performance metrics including social welfare, equality, and win rate, suggests that the AgA approach surpasses the performance of baseline methods.
The contribution of the paper can be summarized as follows: (1) We introduce the differentiable mixed-motive game (DMG) and propose Altruistic Gradient Adjustment (AgA) algorithm to align individual and collective objectives from a gradient perspective. (2) We theoretically prove that careful selection of the alignment weight’s sign can expedite convergence in AgA and effectively avoid unstable fixed points. (3) Visualizations based on a two-player DMG show that our method efficiently aligns individual and collective objectives. Further experiments conducted on mixed-motive testbeds and modified SMAC environments underscore the superior performance of our AgA method in several measurement metrics compared to baselines.
@@ 2 Related Work
Mixed-motive cooperation. Mixed-motive cooperation refers to scenarios where the group’s objectives are sometimes aligned and at other times conflicted (Philip S. Gallo & McClintock, 1965). Recently, there has been a surge in academic interest in the Sequential Social Dilemma (SSD) (Leibo et al., 2017), which expands the concept from its roots in matrix games (Macy & Flache, 2002) to Markov games. Prosocial (Peysakhovich & Lerer, 2018) improves collective performance by blending individual rewards to redraft the agent’s overall utility. While PED-DQN enables agents to incrementally adjust their reward functions for enhanced collaborative action through inter-agent evaluative signal exchanges (Hostallero et al., 2020), Gifting directly reward other agents as part of action space (Lupu & Precup, 2020). Inequity aversion further integrates the concept into Markov games by adding envy (disadvantageous inequality) and guilt (advantageous inequality) rewards to the original individual rewards (Hughes et al., 2018). Simultaneously, Social Value Orientation (SVO) strategy introduces a unique shared rewards-based compensation approach, encouraging behavior modifications in line with interdependence theory (McKee et al., 2020b). The LIO strategy bypasses the need to modify extrinsic rewards by empowering an agent to directly influence its partner’s actions (Yang et al., 2020). Meanwhile, other studies introduce new cooperative mechanisms such as incorporating a reputation model as an intrinsic reward (McKee et al., 2023). The CNM strategy delineates acceptable and unacceptable behaviors using classifiers as social norms, aiding agents to learn advantageous norms for improved returns per agent (Vinitsky et al., 2023). Despite these advances, many current studies lack both cost-efficiency in design and theoretical analysis of alignment and convergence. To address this, our study applies a gradient perspective to align goals and further investigates the learning dynamics of our proposed method.
@@
Gradient-based Methods.  Our proposed AgA is fundamentally a gradient adjustment methodology, making gradient-based optimization methods highly relevant to the context of this paper. Methods based on gradient have been developed to find stationary points, such as the Nash equilibrium or stable fixed points. Optimistic Mirror Descent leverages historical data to extrapolate subsequent gradients (Daskalakis et al., 2018), while Gidel et al. (2020) extends this concept by advocating averaging methodologies and variants of extrapolation techniques. Consensus gradient adjustment, or consensus optimization, is a technique that embeds a consensus agreement term within the gradient to ensure its convergence (Mescheder et al., 2017). Learning with Opponent-Learning Awareness (LOLA) utilizes information from other players to compute one player’s anticipated learning steps (Foerster et al., 2018). Subsequently, Stable Opponent Shaping (SOS) (Letcher et al., 2021) and Consistent Opponent-Learning Awareness (COLA) (Willi et al., 2022) methods have enhanced the LOLA algorithm, targeting convergence assurance and inconsistency elimination, respectively. Symplectic Gradient Adjustment (SGA) alters the update direction towards the stable fixed points based on a novel decomposition of game dynamics (Balduzzi et al., 2018; Letcher et al., 2019). Recently, Learning to Play Games (L2PG) ensures convergence towards a stable fixed point by predicting updates to players’ parameters derived from historical trajectories (Chen et al., 2023). However, these methods, while focusing on zero-sum or general sum games, could potentially act counter to their individual interests, as they may prioritize stability over minimizing personal loss. Our research specifically targets the mixed-motive setting with the aim of reconciling individual and collective objectives.
@@
Figure 1: Trajectories of optimization in a two-player DMG (as delineated in Example 4.1). Fig. 0(a) displays the trajectories over the collective reward landscape - deeper orange equates to higher rewards. Remarkably, only Simul-Co and AgA make successful strides towards the social optimum. Fig. 0(b) and Fig. 0(c) delineate trajectories on the individual reward contour, underscoring Simul-Co’s neglect for Player 1’s interests even though it achieves higher rewards with fewer steps (65 steps versus 100 steps of AgA). Conversely, our AgA optimizes along the summit of Player 1’s reward while also maximizing the collective reward, demonstrating a successful alignment.
3 Preliminaries
Differential Game. The theory of differential games was initially proposed by Isaacs (1965), aiming to expand the scope of sequential game theory to encompass continuous-time scenarios. Through the lens of machine learning, we formalize the differential game, as shown in Definition 3.1.
A differential game could be defined as a tuple $\{{\mathcal{N}},{\bm{w}},{\bm{\ell}}\}$, where ${\mathcal{N}}=\{1,\dots,n\}$ delineates the assembly of players. The parameter set ${\bm{w}}=[{\bm{w}}_{i}]^{n}\in\mathbb{R}^{d}$ is defined, each with ${\bm{w}}_{i}\in\mathbb{R}^{d_{i}}$ and $d=\sum_{i=1}^{n}d_{i}$. Here, =${\bm{\ell}}=\{\ell_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}\}_{i=1}^{n}$ represents the corresponding losses. These losses are assumed to be at least twice differential. Each player $i\in{\mathcal{N}}$ is equipped with a policy, parameterized by ${\bm{w}}_{i}$, aiming to minimize its individual loss $\ell_{i}$.
We rewrite the simultaneous gradient ${\bm{\xi}}({\bm{w}})$ of a differential game as ${\bm{\xi}}({\bm{w}})=(\nabla_{{\bm{w}}_{1}}\ell_{1},\dots,\nabla_{{\bm{w}}_{n}}\ell_{n})\in\mathbb{R}^{d},$ which is the gradient of the losses with respect to the parameters of the respective players. Furthermore, Hessian matrix ${\bm{H}}$ mentioned in a differential game is the Jacobian matrix of the simultaneous gradient.
The learning dynamics of differential game often refers to the process of sequential updates over ${\bm{w}}$. The learning rule for each player is defined as the operator, ${\bm{w}}\leftarrow{\bm{w}}-\gamma{\bm{\xi}},$ where ${\bm{\xi}}$ is a feasible direction, and $\gamma$ is a step size (learning rate) to determine the distance to move for each update.
Gradient Adjustment Optimization. The stable fixed point, a criterion initiated from stability theory, corresponds to the nature of local Nash equilibria (Balduzzi et al., 2018), as defined within the field of game theory. More specifically, a stable fixed point, exhibiting its stability (robustness) to minor perturbations from environments, making it applicable to many real-world scenarios. The concept of stable fixed point prevailing in general-sum and zero-sum games, is defined as follows.
Definition 3.2.
A naive idea to steer the dynamic towards convergence at stable fixed points involves minimizing $\frac{1}{2}\|{\bm{\xi}}({\bm{w}})\|^{2}$. Assuming the Hessian ${\bm{H}}({\bm{w}})$ is invertible, then $\nabla(\frac{1}{2}\|{\bm{\xi}}({\bm{w}})\|^{2})={\bm{H}}^{T}{\bm{\xi}}=0$ holds true if and only if ${\bm{\xi}}=0$. However, it could fail to converge to stable fixed point (Mescheder et al., 2017). Hence, the consensus optimization method has been proposed, incorporating gradient adjustment  (Mescheder et al., 2017), shown as follows:
$\tilde{{\bm{\xi}}}={\bm{\xi}}+\lambda\cdot\nabla\frac{1}{2}\|{\bm{\xi}}({\bm{w}})\|^{2}={\bm{\xi}}+\lambda\cdot{\bm{H}}^{T}{\bm{\xi}}.$ (1)
For simplicity, we will refer to the consensus gradient adjustment as CGA. While CGA proves effective in certain specific scenarios, such as two-player zero-sum games, it unfortunately falls short in general games (Balduzzi et al., 2018). To address this shortage, symplectic gradient adjustment (SGA(Balduzzi et al., 2018; Letcher et al., 2019) is introduced to find the stable fixed point in general sum games, such that $\tilde{{\bm{\xi}}}={\bm{\xi}}+\lambda\cdot{\bm{A}}^{T}{\bm{\xi}}.$ Herein, ${\bm{A}}$ represents the antisymmetric component of the generalized Helmholtz decomposition of ${\bm{H}}$, ${\bm{H}}({\bm{w}})={\bm{S}}({\bm{w}})+{\bm{A}}({\bm{w}})$, where ${\bm{S}}({\bm{w}})$ denotes the symmetric component. The hyperparameter $\lambda$ is determined by the following formula: $\lambda=\operatorname{sign}\left(\frac{1}{d}\langle\bm{\xi},{\bm{H}}^{T}{\bm{\xi}}\rangle\langle\mathbf{A}^{\top}\bm{\xi},{\bm{H}}^{T}{\bm{\xi}}\rangle+\epsilon\right),$ where $\epsilon$ is a small positive number.

### v2

@@ Abstract
Among the research topics in multi-agent learning, mixed-motive cooperation is one of the most prominent challenges, primarily due to the mismatch between individual and collective goals. The cutting-edge research is focused on incorporating domain knowledge into rewards and introducing additional mechanisms to incentivize cooperation. However, these approaches often face shortcomings such as the effort on manual design and the absence of theoretical groundings. To close this gap, we model the mixed-motive game as a differentiable game for the ease of illuminating the learning dynamics towards cooperation. More detailed, we introduce a novel optimization method named Altruistic Gradient Adjustment (AgA) that employs gradient adjustments to progressively align individual and collective objectives. Furthermore, we theoretically prove that AgA effectively attracts gradients to stable fixed points of the collective objective while considering individual interests, and we validate these claims with empirical evidence. We evaluate the effectiveness of our algorithm AgA through benchmark environments for testing mixed-motive collaboration with small-scale agents such as the two-player public good game and the sequential social dilemma games, Cleanup and Harvest, as well as our self-developed large-scale environment in the game StarCraft II.
@@ 1 Introduction
Multi-agent cooperation primarily focuses on learning how to promote collaborative behavior in shared environments. In general, multi-agent cooperation research is categorized into two prominent areas: pure-motive cooperation and mixed-motive cooperation (Du et al., 2023; McKee et al., 2020a). Recent progress in cooperative Multi-Agent Reinforcement Learning (MARL) has been primarily focusing on pure-motive cooperation, also known as common payoff games. This game models situations where each agent’s individual goal fully aligns with the collective objectives (Yu et al., 2022; Wang et al., 2020; Zhong et al., 2023; Li et al., 2024). Nevertheless, mixed-motive cooperation is more widespread across real-world situations. It is usually defined by imperfect alignment between individual and collective rationalities (Rapoport, 1974; McKee et al., 2020a).
@@
Recent studies in mixed-motive cooperative MARL have largely employed hand-crafted designs to promote collaboration, especially with focus on automatically modifying rewards. One popular approach is to align objectives as per existing mechanisms in cooperative games, such as reputation (Anastassacos et al., 2021), norms (Vinitsky et al., 2023), and contracts (Hughes et al., 2020). Another prevalent method leverages intrinsic motivation to align individual and collective objectives, enhancing altruistic collaboration by integrating heuristic knowledge into the incentive function. Conventionally, some works accumulate individual rewards with the group to promote altruistic conduct (Hostallero et al., 2020; Peysakhovich and Lerer, 2018; Apt and Schäfer, 2014; Roesch et al., 2024). Furthermore, some studies derive more sophisticated preference signals from the rewards of other agents (Hughes et al., 2018; McKee et al., 2020a). Additionally, several approaches aim to learn the potential influences of an agent’s actions on others (Yang et al., 2020; Jaques et al., 2019; Lu et al., 2022). Most of these algorithms rely heavily on carefully crafted designs, necessitating significant human expertise and detailed domain knowledge. On the other hand, several studies focus on automatically modifying rewards by learning additional weights to adjust the original objectives (Gemp et al., 2020; Kwon et al., 2023). However, these studies often suffer from a lack of interpretability and insufficient theoretical analysis regarding the alignment of individual and collective objectives.
In this study, we introduce the differentiable mixed-motive game (DMG), an effective framework for analyzing learning dynamics at both individual and collective levels. Furthermore, we derive the Altruistic Gradient Adjustment (AgA) algorithm, which aligns individual and collective objectives by modifying the gradient. We theoretically prove that AgA, with an appropriately chosen sign for the adjustment term, can successfully guide the gradient towards stable fixed points of the collective objective while considering individual interests. Additionally, empirical evidence from optimization trajectory visualizations and ablation studies validates our claims.
We also conduct comprehensive experiments to verify the effectiveness of the proposed AgA algorithm. First, optimization trajectory analysis and ablation studies validate our theoretical conclusions. Next, a series of experiments conducted in various environments demonstrate that the AgA algorithm outperforms related baselines in both gradient adjustment and mixed-motive cooperation areas. In addition to commonly used testbeds like the public goods matrix game and sequential social dilemma games (Cleanup and Harvest) (Leibo et al., 2017), which are limited in terms of agent scale, action space, and task complexity, we introduce a more complex mixed-motive environment called Selfish-MMM2, an adaptation of the MMM2 map from the StarCraft II game (Samvelyan et al., 2019). Selfish-MMM2 offers the following significant improvements over other mixed-motive games: it supports large-scale scenarios with 10 heterogeneous controlled agents facing 11 enemies, features a large action space with a size of $18^{10}$, which vastly exceeds the action spaces in Cleanup $9^{5}$ and Harvest $8^{5}$ that are limited to 5 homogeneous agents. Selfish-MMM2 also has a larger action space and greater task complexity than the 10-player PD testbed (with a size of $2^{10}$) used in D3C (Gemp et al., 2020), which claims to solve the large-scale problem in mixed-motive games.
In summary, the contributions of this paper are as follows: (1) We are the first work to model the mixed-motive game as a differentiable game (to the best of our knowledge) and propose the AgA algorithm to align individual and collective objectives from a gradient perspective. (2) We theoretically prove that, in the neighborhood of fixed points, AgA could pull the gradient toward stable fixed points of the collective objectives and push the gradient away from unstable fixed points. (3) We introduce Selfish-MMM2, a novel large-scale mixed-motive cooperation environment, and conduct comprehensive experiments across multiple settings that verify our theoretical claims and demonstrate the superior performance of the AgA algorithm.
@@ 2 Related Work
Mixed-motive cooperation refers to scenarios where the group’s objectives are sometimes aligned and at other times conflicted (Philip S. Gallo and McClintock, 1965). Recently, there has been a surge in academic interest in the Sequential Social Dilemma (SSD) (Leibo et al., 2017), which expands the concept from its roots in matrix games (Macy and Flache, 2002) to Markov games. Prosocial (Peysakhovich and Lerer, 2018) improves collective performance by blending individual rewards to redraft the agent’s overall utility. While PED-DQN enables agents to incrementally adjust their reward functions for enhanced collaborative action through inter-agent evaluative signal exchanges (Hostallero et al., 2020), Gifting directly reward other agents as part of action space (Lupu and Precup, 2020). Inequity aversion further integrates the concept into Markov games by adding envy (disadvantageous inequality) and guilt (advantageous inequality) rewards to the original individual rewards (Hughes et al., 2018). Simultaneously, Social Value Orientation (SVO) strategy introduces a unique shared rewards-based compensation approach, encouraging behavior modifications in line with interdependence theory (McKee et al., 2020b). The LIO strategy bypasses the need to modify extrinsic rewards by empowering an agent to directly influence its partner’s actions (Yang et al., 2020). Meanwhile, other studies introduce new cooperative mechanisms such as incorporating a reputation model (McKee et al., 2023), social norms (Vinitsky et al., 2023). Recently, Roesch et al. (2024) proposed a selfishness level method that incorporates social welfare into individual rewards to enhance altruistic cooperation. Additionally, several works have explored automatically modifying rewards online. D3C (Gemp et al., 2020) learns to mix rewards to improve efficiency in a Nash equilibrium. Kwon et al. (2023) extended D3C by addressing the problem of automatically modifying individual agent objectives to optimize a desired global objective. Despite these advances, many current studies lack both cost efficiency in design and theoretical analysis of alignment and convergence. To address this, our study applies a gradient perspective to align goals and further investigates the learning dynamics of our proposed method.
Gradient-based Methods.
@@
Our proposed AgA is fundamentally a gradient adjustment methodology, making gradient-based optimization methods highly relevant to the context of this paper. Methods based on gradient have been developed to find stationary points, such as the Nash equilibrium or stable fixed points. Optimistic Mirror Descent leverages historical data to extrapolate subsequent gradients Daskalakis et al. (2018), while Gidel et al. (2020) extends this concept by advocating averaging methodologies and variants of extrapolation techniques. Consensus gradient adjustment, or consensus optimization, is a technique that embeds a consensus agreement term within the gradient to ensure its convergence (Mescheder et al., 2017). Learning with Opponent-Learning Awareness (LOLA) utilizes information from other players to compute one player’s anticipated learning steps (Foerster et al., 2018). Subsequently, Stable Opponent Shaping (SOS) (Letcher et al., 2021) and Consistent Opponent-Learning Awareness (COLA) (Willi et al., 2022) methods have enhanced the LOLA algorithm, targeting convergence assurance and inconsistency elimination, respectively. Symplectic Gradient Adjustment (SGA) alters the update direction towards the stable fixed points based on a novel decomposition of game dynamics (Balduzzi et al., 2018; Letcher et al., 2019). Recently, Learning to Play Games (L2PG) ensures convergence towards a stable fixed point by predicting updates to players’ parameters derived from historical trajectories (Chen et al., 2023). However, these methods, while focusing on zero-sum or general sum games, could potentially act counter to their individual interests, as they may prioritize stability over minimizing personal loss. Our research specifically targets mixed-motive setting with the aim of reconciling individual and collective objectives.
3 Preliminaries
3.1 Differential Game
The theory of differential games was initially proposed by Isaacs (1965), aiming to expand the scope of sequential game theory to encompass continuous-time scenarios. Through the lens of machine learning, we formalize the differential game, as shown in Definition 3.1.
A differential game could be defined as a tuple $\{{\mathcal{N}},{\bm{w}},{\bm{\ell}}\}$, where ${\mathcal{N}}=\{1,\dots,n\}$ is the set of players. The parameter set ${\bm{w}}=[{\bm{w}}_{i}]^{n}\in\mathbb{R}^{d}$ is defined, each with ${\bm{w}}_{i}\in\mathbb{R}^{d_{i}}$ and $d=\sum_{i=1}^{n}d_{i}$. Here, =${\bm{\ell}}=\{\ell_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}\}_{i=1}^{n}$ represents the corresponding losses. These losses are assumed to be at least twice differential. Each player $i\in{\mathcal{N}}$ is equipped with a policy, parameterized by ${\bm{w}}_{i}$, aiming to minimize its loss $\ell_{i}$.
We write the simultaneous gradient ${\bm{\xi}}({\bm{w}})$ of a differential game as ${\bm{\xi}}({\bm{w}})=(\nabla_{{\bm{w}}_{1}}\ell_{1},\dots,\nabla_{{\bm{w}}_{n}}\ell_{n})\in\mathbb{R}^{d},$ which is the gradient of the losses with respect to the parameters of the respective players. Furthermore, Hessian matrix ${\bm{H}}$ mentioned in a differential game is the Jacobian matrix of the simultaneous gradient.
The learning dynamics of differential game often refers to the process of sequential updates over ${\bm{w}}$. The learning rule for each player is defined as the operator, ${\bm{w}}\leftarrow{\bm{w}}-\gamma{\bm{\xi}},$ where $\gamma$ is a step size (learning rate) to determine the distance to move for each update.
3.2 Gradient Adjustment Optimization
The stable fixed point, a criterion initiated from stability theory, exhibits its stability (robustness) to minor perturbations of environments, making it applicable to many real-world scenarios.
Definition 3.2.
A naive idea to steer the dynamic towards convergence at fixed points involves minimizing $\frac{1}{2}\|{\bm{\xi}}({\bm{w}})\|^{2}$. Assuming the Hessian ${\bm{H}}({\bm{w}})$ is invertible, then $\nabla(\frac{1}{2}\|{\bm{\xi}}({\bm{w}})\|^{2})={\bm{H}}^{T}{\bm{\xi}}=0$ holds true if and only if ${\bm{\xi}}=0$. However, it could converge to unstable fixed points (Mescheder et al., 2017). Hence, the consensus optimization method has been proposed, incorporating gradient adjustment  (Mescheder et al., 2017), shown as follows: $\tilde{{\bm{\xi}}}={\bm{\xi}}+\lambda\cdot\nabla\frac{1}{2}\|{\bm{\xi}}({\bm{w}})\|^{2}={\bm{\xi}}+\lambda\cdot{\bm{H}}^{T}{\bm{\xi}}.$ For simplicity, we will refer to the consensus gradient adjustment as CGA. While CGA proves effective in certain specific scenarios, such as two-player zero-sum games, it unfortunately falls short in general games (Balduzzi et al., 2018). To address this shortage, symplectic gradient adjustment (SGA(Balduzzi et al., 2018; Letcher et al., 2019) is introduced to find the stable fixed point in general sum games, such that $\tilde{{\bm{\xi}}}={\bm{\xi}}+\lambda\cdot{\bm{A}}^{T}{\bm{\xi}}.$ Herein, ${\bm{A}}$ represents the antisymmetric component of the generalized Helmholtz decomposition of ${\bm{H}}$, ${\bm{H}}({\bm{w}})={\bm{S}}({\bm{w}})+{\bm{A}}({\bm{w}})$, where ${\bm{S}}({\bm{w}})$ denotes the symmetric component.
@@ 3.2 Gradient Adjustment Optimization
Figure 1: Trajectories of optimization in a two-player DMG (as delineated in Example 4.1). Fig.1(a) displays the trajectories over the collective reward landscape - deeper orange equates to higher rewards. Remarkably, only Simul-Co and AgA make successful strides towards the social optimum. Fig.1(b) and Fig. 1(c) delineate trajectories on the individual reward contour, underscoring Simul-Co’s neglect for Player 1’s interests as it navigates through the crests and troughs of its reward. Conversely, our AgA optimizes along the summit of Player 1’s reward while also maximizing the collective reward, demonstrating successful alignment.